Circular No. 8859

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 O2 = P/1998 U2 (MUELLER) L. Buzzi and F. Luppi, Varese, Italy, report the recovery of comet

P/1998 U2 (cf. IAUC 7035, 7036), being slightly diffuse on CCD images obtained with a 0.60-m reflector, with confirming CCD images by P. Birtwhistle (Great Shefford, Berkshire, U.K., 0.40-m reflector).

2007 UT α_{2000} δ_{2000} Mag. Observer
July 26.09058 3^{h}48^{m}28.^{s}53 +21^{o}41′05′′.3 19.5 Buzzi
26.10107 3 48 29.86 +21 41 09.8  _{”}
27.10651 3 50 39.81 +21 48 37.6  Birtwhistle
27.11524 3 50 40.92 +21 48 43.8 19.5  _{”}

The indicated correction to the predictions in the 2007 Comet Handbook and on MPC 51823 is ΔT = +0.3 day. The following orbital elements by B. G. Marsden are from 153 observations, 1998-2007 (mean residual 0′′.6):

Epoch = 1998 Nov. 3.0 TT

T = 1998 Oct. 20.75954 TT  ω = 49.^{o}53569
e = 0.5215660  Ω = 336.22405 }2000.0
q = 2.0253092 AU  i = 2.19329

a = 4.2332052 AU n^{o} = 0.11316187 P = 8.71 years Epoch = 2007 June 29.0 TT

T = 2007 July 8.20087 TT  ω = 49.^{o}73116
e = 0.5206516  Ω = 336.11998 }2000.0
q = 2.0319152 AU  i = 2.18993

a = 4.2389107 AU n^{o} = 0.11293347 P = 8.73 years

R CORONAE BOREALIS E. Broens, Vereniging Voor Sterrenkunde, Belgium, reports that this

variable has begun to fade slowly for the first time since Feb.–May 2003 (cf. IAUC 8077), when it faded over ∼ 20 days from a visual mag ∼ 6.1 to minimum at mag ∼ 12.7. Visual magnitude estimates (non-Belgian data sent directly to the Central Bureau): June 21.94 UT, 6.0 (E. Muyllaert, Oostende, Belgium); July 7.93, 6.3 (Muyllaert); 13.85, 6.5 (F. Van Loo, Genk, Belgium); 16.878, 7.1 (S. Baroni, Milan, Italy); 17.837 7.2 (S. Foglia, Novara, Italy); 18.909, 7.0 (K. Hornoch, Vohancice, Czech Republic); 19.896, 7.0 (Baroni); 20.96, 7.1 (E. Broens, Mol, Belgium); 22.881, 7.3 (Hornoch); 24.862, 7.6 (Foglia); 24.920, 7.4 (Hornoch).

2007 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green